A street in manila composing of two lanes each 14ft wide runs through a semi circular tunnel with a radius of 16ft. What is the height of the tunnel at the edge of each lane?
Answer by rothauserc(4718)

for a circle, we know x^2 +y^2 = r^2, where r is the radius of the circle
:
therefore, for the semicircle we have y = square root(r^2 -x^2), where the square root is positive for the semi circular tunnel
:
let the center for the semicircle tunnel be (0,0), then
:
the outside edges of the two lanes are -14 and 14
:
Note that y is the height of the tunnel
:
y = square root(256-x^2)
:
y = square root(256-196) = 7.746
:
the tunnel is 7.746 feet high at each edge
: